Bill is 4 less than twice Alex's age. Together they are 65 years old. How old is Bill?
Answer by mananth(16946)

Alex = x
Bill = 2x-4
...
x+2x-4 = 65
3x-4=65
+4
3x= 69
x= 23 years Alex age
...
Bill = 65-23 = 42 years